Finding the slope and y-intercept for each line that has a slope and y-intercept. The problem is x+2y=3
Answer by Nate(3500)

Put the equation in slope intercept form for easy access to information.
x + 2y = 3
2y = -x + 3
y = -x/2 + 3/2
slope-intercept form: y = mx + b where m is the slope and b is the y-intercept
-1/2 = slope
3/2 = y-intercept